OpenStudy (anonymous):

Lindy is planning to build a model of a train using a scale where 2 inches represents 25 feet. If the train is 60 feet long, what is the length in inches that Lindy should build the

OpenStudy (anonymous):

\[\frac{ 2 inch }{ 25 feet }=\frac{ x inch }{ 60 feet }\] \[\frac{ 2x60 }{ 25 }=x=4.8 inch ..ans\]

OpenStudy (anonymous):

thank u!

OpenStudy (anonymous):

yw, good luck